A point moves such that the sum of the squares of its distance from the sides of a square of side unity is equal to 9. Show that the locus is a circle whose centre coincides with the centre of the square. Find also its radius

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Debadutta

Let us take a simple square having four vertices at (.5,.5),(-.5,.5),(.5,-.5),(-.5,-.5),

Point is (x1,y1)

So according to question

(x1+.5)^2+(y1+.5)^2+(x1+.5)^2+(y1-.5)^2+(x1-.5)^2+(y1+.5)^2+(x1-.5)^2+(y1-.5)^2=9

4x12+4y12=9-8*.52=7

So we can see locus is a circle whose centre coincides with the centre of the square and radius of circle is square root of 7/4.

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..